# positioning

### v1

@@ Abstract.
Fix distinct primes $p$ and $q$ and let $E$ be an elliptic curve defined over a number field $K$. The $(p,q)$-entanglement type of $E$ over $K$ is the isomorphism class of the group $\operatorname{Gal}(K(E[p])\cap K(E[q])/K)$. This size of this group measures the extent to which the image of the mod $pq$ Galois representation attached to $E$ fails to be a direct product of the mod $p$ and mod $q$ images. In this article, we study how the $(p,q)$-entanglement group varies over different base fields. We prove that for each prime $\ell$ dividing the greatest common divisor of the size of the mod $p$ and $q$ images, there are infinitely many fields $L/K$ such that the entanglement over $L$ is cyclic of order $\ell$. We also classify all possible $(2,q)$-entanglement types that can occur as the base field $L$ varies.
@@ 1. Introduction
whose image is well-defined up to conjugation. A classical problem, often referred to as Mazur’s ‘Program B’ [Maz77], is to determine all possible images of this representation. When $n=p$ is prime and $E$ does not have complex multiplication, a well-known theorem of Serre [Ser72] asserts that $\rho_{E,p}$ is surjective for all but finitely many exceptional primes $p$. Recent work of Zywina [Zyw15] classifies the possible images of $\rho_{E,p}$ for rational elliptic curves at exceptional primes.
@@
{tikzpicture}
{tikzpicture}
@@
If both the mod $p$ and $q$ representations of $E/K$ are surjective (true for all but finitely many $(p,q)$ by [Ser72]) then
@@
The size of $d(E,p,q)$ can be drastically smaller when either $p$ or $q$ is exceptional, or when $E$ has complex multiplication, and in these cases the divisibility (2) forces $\operatorname{Ent}_{p,q}(E/K)$ to also shrink in size. We remark that, for any given elliptic curve $E/\mathbb{Q}$, the number $d(E,p,q)$ is explicitly computable using [Zyw15].
@@ 1.1. Related questions and results
There has been much work in recent years on classifying all elliptic curves defined over a fixed base field $K$ (often $K=\mathbb{Q}$) having some prescribed entanglement properties. See, for example, [BJ16], [Mor19], [DM20], [DLM21], and [JM22], where variations of the following question are considered.
@@
This type of question is amenable to study using the theory of modular curves, where (roughly speaking) one constructs a modular curve whose $K$-rational points correspond to elliptic curves with the desired entanglement properties. For example, by considering certain genus 0 and genus 1 (positive rank) modular curves, it is shown in [DM20] that the only unexplained (see Definition 4.5 of op. cit.) entanglement types $T$ which can occur for infinitely many $\bar{\mathbb{Q}}$-isomorphic elliptic curves are $\mathbb{Z}/2\mathbb{Z}$, $\mathbb{Z}/3\mathbb{Z}$, and $S_{3}$, and that they can occur only at the pairs $(p,q)$ in the sets $\{(2,3),(2,5),(2,7),(2,13),(3,5)\}$, $\{(2,3),(2,5),(2,7)\}$, and $\{(2,3)\}$, respectively.
@@
    For fixed $(K,p,q)$, does there always exist an elliptic curve $E/K$ having nontrivial $(p,q)$-entanglement over $K$? When $K=\mathbb{Q}$ and $p=2$, the answer to this question is ‘yes’: one can force explained (see [DM20]) entanglement using the Weil pairing. Specifically, any elliptic curve $E/\mathbb{Q}$ with $j$-invariant $1728\pm qt^{2},$ where $\pm$ is the sign of $(-1)^{\frac{q-1}{2}}$ and $t\in\mathbb{Q}^{\times}$, will have $(2,q)$-entanglement: the $j$-invariant forces $\Delta_{E}\equiv\pm q\ \mathrm{mod}\ \mathbb{Q}^{\times 2}$, thus $\mathbb{Q}(\sqrt{\Delta_{E}})=\mathbb{Q}(\sqrt{\pm q})$ is contained in both $\mathbb{Q}(E[2])$ and $\mathbb{Q}(\zeta_{q})\subseteq\mathbb{Q}(E[q])$.
@@
1.2. Acknowledgements
The authors would like to thank Harris B. Daniels for many inspiring conversations and crucial suggestions throughout the preparation of this article. We are also grateful to Jeffrey Hatley for several helpful comments and corrections.
@@ 2.1. Group-theoretic formulation
Fix a basis for $E[pq]$ and let $\rho_{E,pq}:G_{K}\rightarrow\operatorname{GL}_{2}(\mathbb{Z}/pq\mathbb{Z})$ denote the Galois representation corresponding to the action of $G_{K}$ on $E[pq]$. Following [DM20], the group $\mathcal{G}:=\operatorname{im}\rho_{E,pq}$ is said to represent a $(p,q)$-entanglement if
@@ 2.2. Stable and decreasing entanglement
Follows from the fact that if $L\subseteq K_{p,q}$ then $L(E[p])=K(E[p])$ and $L(E[q])=K(E[q])$. ∎
@@
for some $t\in\mathbb{Q}\backslash\{0,\frac{1}{2}\}$. (For instance, taking $t=\frac{2}{3}$ gives the elliptic curve 300a1.) Then it follows from [BJ16, Remark 1.5] and [DM20, §8.1], respectively, that
@@
The relation between $T_{p,q}(E/L)$ and $T_{p,q}(E/K)$ for arbitrary $L/K$ is illustrated in the following field diagram:
{tikzpicture}
@@
Then $LK_{1}$ and $LK_{2}$ are linearly disjoint over $L(K_{1}\cap K_{2})$ if and only if $[LK_{1}K_{2}:L(K_{1}\cap K_{2})]=a_{1}a_{2}$. It therefore suffices to show that $b=1$ if and only if $[LK_{1}K_{2}:L(K_{1}\cap K_{2})]=a_{1}a_{2}$, which follows easily from the diagram below, where the dashed lines indicate extensions of the same degree.
{tikzpicture}
@@
From the Lemma 2.5, we know that $K_{p,q}L=L_{p,q}$. The result now follows from the field diagram in Figure 1. ∎
@@
The following lemma allows one to determine the size of the entanglement along subfields of $K(E[pq])$.
@@ 3. Entangleable fields
The coprimality hypothesis in Proposition 3.4 is essentially the only obstruction to fields being entangleable (see Proposition 3.11 below). In order to see this, we recast field entangleability in group-theoretic terms.
Let $F/K$ be any finite Galois extension containing both $F_{1}$ and $F_{2}$ and write $G=\operatorname{Gal}(F/K)$, $G_{1}=\operatorname{Gal}(F/F_{1})$, and $G_{2}=\operatorname{Gal}(F/F_{2})$.
@@
$F_{1}$ and $F_{2}$ are entangleble by a subfield $L\subseteq F$ if and only if there exists a subgroup $H\leq G$ such that $H\cap G_{1}=H\cap G_{2}\subsetneq H$. If this is the case, then $L$ and $H$ are related by the correspondence
$L=F^{H}\longleftrightarrow H=\operatorname{Gal}(F/L),$
and $\operatorname{Gal}(LF_{i}/L)\cong H/(H\cap G_{i})$.
@@
Let $D_{1}$ and $D_{2}$ be squarefree integers. Let $\sigma$ and $\tau$ be generators of $G=\operatorname{Gal}(\mathbb{Q}(\sqrt{D_{1}},\sqrt{D_{2}})/\mathbb{Q})$ that restrict to the nontrivial elements of $G_{1}=\operatorname{Gal}(\mathbb{Q}(\sqrt{D_{1}})/\mathbb{Q})$ and $G_{2}=\operatorname{Gal}(\mathbb{Q}(\sqrt{D_{2}})/\mathbb{Q})$, respectively. Then the subgroup $H\leq G$ generated by $\sigma\tau$ trivially intersects $G_{1}$ and $G_{2}$, hence $\mathbb{Q}(\sqrt{D_{1}})$ and $\mathbb{Q}(\sqrt{D_{2}})$ are entangleable by $\mathbb{Q}(\sqrt{D_{1}},\sqrt{D_{2}})^{H}=\mathbb{Q}(\sqrt{D_{1}D_{2}})$.
@@ 3.1. Criteria for field entangleability
under the mapping $g\mapsto(g|_{F_{1}},g|_{F_{2}})$.
Let $\ell$ be a prime divisor of $d$. By Cauchy’s theorem, there are elements $g_{i}\in\operatorname{Gal}(F_{i}/F_{1}\cap F_{2})$ of order $\ell$. Let $H$ be the preimage in $G$ of the subgroup $\langle(g_{1},g_{2})\rangle$ of the fibre product above. (Note that $(g_{1},g_{2})$ is indeed an element of the fibre product.) Then $H$ has order $\ell$ and $H\cap G_{i}$ is trivial. (Why? Let $g\in G$ be the element mapping to $(g_{1},g_{2})$, so $H=\langle g\rangle$. If $g\in G_{i}=\operatorname{Gal}(F_{1}F_{2}/F_{i})$ then $g|_{F_{i}}=1=g_{i}$, a contradiction.) It follows from Lemma 3.5 that $F_{1}$ and $F_{2}$ are entangleable by $L=F^{H}$. ∎
@@ 4.1. Cyclic entanglements
First, we claim that it suffices to prove that there exist subfields $F_{1},F_{2}$, and $L$ of $K(E[pq])$ such that $F_{1}$ and $F_{2}$ are entangleable by $L$ and $T_{p,q}(E/L)=\mathbb{Z}/\ell\mathbb{Z}.$ Assuming that this is true, let $L^{\prime}/K$ be any field for which $LL^{\prime}\cap F_{1}=L\cap F_{1}$. By Proposition 3.14, there are infinitely many such $L^{\prime}$. By Proposition 3.13 (and its proof), we know that $F_{1}$ and $F_{2}$ are entangleable by $LL^{\prime}$ and
(6) $\operatorname{Gal}(LL^{\prime}F_{1}/LL^{\prime})\cong\operatorname{Gal}(LF_{1}/L).$
Now, since $F_{1}$ and $F_{2}$ are entangleable by $L$, Theorem 4.1 implies
@@
and since $F_{1}$ and $F_{2}$ are also entangleable by $LL^{\prime}$, again Theorem 4.1 implies
@@ 5. Classification of $(2,q)$-entanglement types
By Lemma 3.5 (with $F=\mathbb{Q}(E[2q])$, $F_{1}=\mathbb{Q}(E[2])$ and $F_{2}=\mathbb{Q}(E[q])$) and Proposition 4.1, it suffices to show that $\operatorname{GL}_{2}(\mathbb{Z}/2\mathbb{Z})\times\operatorname{GL}_{2}(\mathbb{Z}/q\mathbb{Z})$ contains a subgroup $H$ isomorphic to $S_{3}$ such that both $H\cap\operatorname{GL}_{2}(\mathbb{Z}/2\mathbb{Z})\times 1$ and $H\cap 1\times\operatorname{GL}_{2}(\mathbb{Z}/q\mathbb{Z})$ are trivial. This fact follows from Lemma 5.2 below. ∎

### v2

# Entanglement of elliptic curves upon base extension

@@ Abstract.
Fix distinct primes $p$ and $q$ and let $E$ be an elliptic curve defined over a number field $K$. The $(p,q)$-entanglement type of $E$ over $K$ is the isomorphism class of the group $\operatorname{Gal}(K(E[p])\cap K(E[q])/K)$. The size of this group measures the extent to which the image of the mod $pq$ Galois representation attached to $E$ fails to be a direct product of the mod $p$ and mod $q$ images. In this article, we study how the $(p,q)$-entanglement group varies over different base fields. We prove that for each prime $\ell$ dividing the greatest common divisor of the size of the mod $p$ and $q$ images, there are infinitely many fields $L/K$ such that the entanglement over $L$ is cyclic of order $\ell$. We also classify all possible $(2,q)$-entanglement types that can occur as the base field $L$ varies.
@@ 1. Introduction
whose image is well-defined up to conjugation. A classical problem, often referred to as Mazur’s ‘Program B’ [6], is to determine all possible images of this representation. When $n=p$ is prime and $E$ does not have complex multiplication, a well-known theorem of Serre [8] asserts that $\rho_{E,p}$ is surjective for all but finitely many exceptional primes $p$. Recent work of Zywina [9] classifies the possible images of $\rho_{E,p}$ for rational elliptic curves at exceptional primes.
@@
$\mathbb{Q}(E[2])$$\mathbb{Q}(E[3])$$\mathbb{Q}(\sqrt{-21})$$\mathbb{Q}(\sqrt{21})$$\mathbb{Q}(\sqrt{-7})$$\mathbb{Q}(\sqrt{-3})$$\mathbb{Q}$
$L(E[2])$$L(E[3])$$L(\sqrt{-21})=L(\sqrt{21})$$L(\sqrt{-7})$$L(\sqrt{-3})$$L$
@@
If both the mod $p$ and $q$ representations of $E/K$ are surjective (true for all but finitely many $(p,q)$ by [8]) then
@@
The size of $d(E,p,q)$ can be drastically smaller when either $p$ or $q$ is exceptional, or when $E$ has complex multiplication, and in these cases the divisibility (2) forces $\operatorname{Ent}_{p,q}(E/K)$ to also shrink in size. We remark that, for any given elliptic curve $E/\mathbb{Q}$, the number $d(E,p,q)$ is explicitly computable using [9].
@@ 1.1. Related questions and results
There has been much work in recent years on classifying all elliptic curves defined over a fixed base field $K$ (often $K=\mathbb{Q}$) having some prescribed entanglement properties. See, for example, [2], [7], [3], [4], and [5], where variations of the following question are considered.
@@
This type of question is amenable to study using the theory of modular curves, where (roughly speaking) one constructs a modular curve whose $K$-rational points correspond to elliptic curves with the desired entanglement properties. For example, by considering certain genus 0 and genus 1 (positive rank) modular curves, it is shown in [3] that the only unexplained (see Definition 4.5 of op. cit.) entanglement types $T$ which can occur for infinitely many $\bar{\mathbb{Q}}$-isomorphic elliptic curves are $\mathbb{Z}/2\mathbb{Z}$, $\mathbb{Z}/3\mathbb{Z}$, and $S_{3}$, and that they can occur only at the pairs $(p,q)$ in the sets $\{(2,3),(2,5),(2,7),(2,13),(3,5)\}$, $\{(2,3),(2,5),(2,7)\}$, and $\{(2,3)\}$, respectively.
@@
    For fixed $(K,p,q)$, does there always exist an elliptic curve $E/K$ having nontrivial $(p,q)$-entanglement over $K$? When $K=\mathbb{Q}$ and $p=2$, the answer to this question is ‘yes’: one can force explained (see [3]) entanglement using the Weil pairing. Specifically, any elliptic curve $E/\mathbb{Q}$ with $j$-invariant $1728\pm qt^{2},$ where $\pm$ is the sign of $(-1)^{\frac{q-1}{2}}$ and $t\in\mathbb{Q}^{\times}$, will have $(2,q)$-entanglement: the $j$-invariant forces $\Delta_{E}\equiv\pm q\ \mathrm{mod}\ \mathbb{Q}^{\times 2}$, thus $\mathbb{Q}(\sqrt{\Delta_{E}})=\mathbb{Q}(\sqrt{\pm q})$ is contained in both $\mathbb{Q}(E[2])$ and $\mathbb{Q}(\zeta_{q})\subseteq\mathbb{Q}(E[q])$.
@@
1.2. Acknowledgments
The authors thank Harris B. Daniels for many inspiring conversations and crucial suggestions throughout the preparation of this article. We are also grateful to Jeffrey Hatley for several helpful comments and corrections.
@@ 2.1. Group-theoretic formulation
Fix a basis for $E[pq]$ and let $\rho_{E,pq}:G_{K}\rightarrow\operatorname{GL}_{2}(\mathbb{Z}/pq\mathbb{Z})$ denote the Galois representation corresponding to the action of $G_{K}$ on $E[pq]$. Following [3], the group $\mathcal{G}:=\operatorname{im}\rho_{E,pq}$ is said to represent a $(p,q)$-entanglement if
@@ 2.2. Stable and decreasing entanglement
This follows from the fact that if $L\subseteq K_{p,q}$ then $L(E[p])=K(E[p])$ and $L(E[q])=K(E[q])$. ∎
@@
for some $t\in\mathbb{Q}\backslash\{0,\frac{1}{2}\}$. (For instance, taking $t=\frac{2}{3}$ gives the elliptic curve 300a1.) Then it follows from [2, Remark 1.5] and [3, §8.1], respectively, that
@@
The relation between $T_{p,q}(E/L)$ and $T_{p,q}(E/K)$ for arbitrary $L/K$ is illustrated in the following diagram.
[…]
@@
Then $LK_{1}$ and $LK_{2}$ are linearly disjoint over $L(K_{1}\cap K_{2})$ if and only if $[LK_{1}K_{2}:L(K_{1}\cap K_{2})]=a_{1}a_{2}$. It therefore suffices to show that $b=1$ if and only if $[LK_{1}K_{2}:L(K_{1}\cap K_{2})]=a_{1}a_{2}$, which follows from the diagram below, where the dashed lines indicate extensions of the same degree.
[…]
@@
From Lemma 2.5, we know that $K_{p,q}L=L_{p,q}$. The result now follows from the field diagram in Figure 1. ∎
@@
The following proposition allows one to determine the size of the entanglement along subfields of $K(E[pq])$.
@@ 3. Entangleable Fields
The coprimality hypothesis in Proposition 3.4 is essentially the only obstruction to fields being entangleable (see Proposition 3.11 below). In order to see this, we recast field entangleability in group-theoretic terms. Let $F/K$ be any finite Galois extension containing both $F_{1}$ and $F_{2}$ and write $G=\operatorname{Gal}(F/K)$, $G_{1}=\operatorname{Gal}(F/F_{1})$, and $G_{2}=\operatorname{Gal}(F/F_{2})$.
@@
$F_{1}$ and $F_{2}$ are entangleable by a subfield $L\subseteq F$ if and only if there exists a subgroup $H\leq G$ such that $H\cap G_{1}=H\cap G_{2}\subsetneq H$. If this is the case, then $L$ and $H$ are related by the correspondence
$L=F^{H}\longleftrightarrow H=\operatorname{Gal}(F/L).$
Furthermore, if $L/K$ or $F_{i}/K$ is Galois then $\operatorname{Gal}(LF_{i}/L)\cong H/(H\cap G_{i})$.
@@
Let $D_{1}$ and $D_{2}$ be squarefree integers. Let $\sigma$ and $\tau$ be generators of $G=\operatorname{Gal}(\mathbb{Q}(\sqrt{D_{1}},\sqrt{D_{2}})/\mathbb{Q})$ that restrict to the nontrivial elements of $\operatorname{Gal}(\mathbb{Q}(\sqrt{D_{1}})/\mathbb{Q})$ and $\operatorname{Gal}(\mathbb{Q}(\sqrt{D_{2}})/\mathbb{Q})$, respectively. Then the subgroup $H\leq G$ generated by $\sigma\tau$ trivially intersects $G_{i}=\operatorname{Gal}(\mathbb{Q}(\sqrt{D_{1}},\sqrt{D_{2}})/\mathbb{Q}(\sqrt{D_{i}}))$, hence $\mathbb{Q}(\sqrt{D_{1}})$ and $\mathbb{Q}(\sqrt{D_{2}})$ are entangleable by $\mathbb{Q}(\sqrt{D_{1}},\sqrt{D_{2}})^{H}=\mathbb{Q}(\sqrt{D_{1}D_{2}})$.
@@ 3.1. Criteria for field entangleability
under the mapping $g\mapsto(g|_{F_{1}},g|_{F_{2}})$. Let $\ell$ be a prime divisor of $d$. By Cauchy’s theorem, there are elements $g_{i}\in\operatorname{Gal}(F_{i}/F_{1}\cap F_{2})$ of order $\ell$. Let $H$ be the preimage in $G$ of the subgroup $\langle(g_{1},g_{2})\rangle$ of the fibre product above. (Note that $(g_{1},g_{2})$ is indeed an element of the fibre product.) Then $H$ has order $\ell$ and $H\cap G_{i}$ is trivial. To see this, let $g\in G$ be the element mapping to $(g_{1},g_{2})$ (thus $H=\langle g\rangle$) and observe that if $g\in G_{i}=\operatorname{Gal}(F_{1}F_{2}/F_{i})$ then $g|_{F_{i}}=1=g_{i}$, a contradiction. It follows from Lemma 3.5 that $F_{1}$ and $F_{2}$ are entangleable by $L=F^{H}$. ∎
@@ 4.1. Cyclic entanglements
First, we claim that it suffices to prove that there exist subfields $F_{1},F_{2}$, and $L$ of $K(E[pq])$ such that $F_{1}/K$ is Galois, $F_{1}$ and $F_{2}$ are entangleable by $L$, and $T_{p,q}(E/L)=\mathbb{Z}/\ell\mathbb{Z}.$ Assuming that this is true, let $L^{\prime}/K$ be any field for which $LL^{\prime}\cap F_{1}=L\cap F_{1}$. By Proposition 3.14, there are infinitely many such $L^{\prime}$. By Proposition 3.13, we know that $F_{1}$ and $F_{2}$ are entangleable by $LL^{\prime}$ and
(6) $[LL^{\prime}F_{1}:LL^{\prime}]=[LF_{1}:L].$
Now, since $F_{1}$ and $F_{2}$ are entangleable by $L$, Proposition 4.1.(1) implies
@@
and since $F_{1}$ and $F_{2}$ are also entangleable by $LL^{\prime}$, again Proposition 4.1.(1) implies
@@ 5. Classification of $(2,q)$-Entanglement Types
By Lemma 3.5 (with $F=\mathbb{Q}(E[2q])$, $F_{1}=\mathbb{Q}(E[2])$, and $F_{2}=\mathbb{Q}(E[q])$) and Proposition 4.1, it suffices to show that $\operatorname{GL}_{2}(\mathbb{Z}/2\mathbb{Z})\times\operatorname{GL}_{2}(\mathbb{Z}/q\mathbb{Z})$ contains a subgroup $H$ isomorphic to $S_{3}$ such that both $H\cap\operatorname{GL}_{2}(\mathbb{Z}/2\mathbb{Z})\times 1$ and $H\cap 1\times\operatorname{GL}_{2}(\mathbb{Z}/q\mathbb{Z})$ are trivial. This fact follows from Lemma 5.2 below. ∎